OpenStudy (aroub):

Can someone please help me in parametric curves? How do you pick the additional points? Or they could be literally any points? & how many do you normally choose?

OpenStudy (amistre64):

a parametric curve is simply defined for each part individually ... would need more context to determine what you are actually asking about tho

OpenStudy (aroub):

Draw the graph given parametrically by: x=x(t)=t^2-4 y=y(t)=t^3-4t

OpenStudy (aroub):

You have to do few steps in order to graph it, and additional points is one of them, but I don't know if it's that important? Can you can graph it without the additional points?

OpenStudy (amistre64):

x=t^2-4 y=t^3-4t t(t^2-4) looks like y = tx

OpenStudy (amistre64):

if you can convert it into a cartesian equation ... it may be easier to see. other than that, you simply make a table of txy values and plot them

OpenStudy (amistre64):

x=t^2-4 x+4 = t^2 +-sqrt(x+4) = t y = x sqrt(x+4) or y = -x sqrt(x+4) http://www.wolframalpha.com/input/?i=y%3Dx+sqrt%28x%2B4%29%2Cy%3D-x+sqrt%28x%2B4%29

OpenStudy (amistre64):

does that make sense?

OpenStudy (aroub):

Oh, nice! I didn't see that. Haha, thanks :)

OpenStudy (amistre64):

good luck :)